OpenStudy (anonymous):

In testing a new drug, researchers found that 5% of all patients using it will have a mild side effect. A random sample of 12 patients using the drug is selected. Find the probability that: (A) exactly two will have this mild side effect (B) at least two will have this mild side effect.

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

as far as the formula goes i'm figuring its n = 12 and um, p = .05

OpenStudy (anonymous):

so i'm under the assumption that i'm working with 12C2(0.05)^2 x .95 or something along those line but i'm not sure how to draw the formula up

OpenStudy (skullpatrol):

5% of 12 patients is 0.6

OpenStudy (anonymous):

Its A. I think :) Binomial Problem with n = 12 ; p = 0.05 P(x=3) = 0.05^3 = 0.125 -------------------------------------------------- (B) at least two will have this mild side effect. P(2<=x<=12) = 1 - P(0<= x <=1) = 0.1184